integration using trigonometric identetiees cos2xsinx

Grade:12

1 Answers

Diksha Chandna
7 years ago
Yaa here`s the answer to your query..Integration(sinx.cosxdx)By using the identity sinAcosB = 1/2[sin(A-B)+sin(A+B)]So, Integration(sinx.cosxdx)We get 1/2int [sin(-x) + sin3x]= 1/2 int[-sinx + sin3x]=1/2[int(-sinx) + int(sin3x)]=1/2(cosx -1/3cos3x) +C